# DriveMLM: Aligning Multi-Modal Large Language Models with Behavioral Planning States for Autonomous Driving

### v2

@@ Abstract
Large language models (LLMs) have opened up new possibilities for intelligent agents, endowing them with human-like thinking and cognitive abilities. In this work, we delve into the potential of large language models (LLMs) in autonomous driving (AD). We introduce DriveMLM, an LLM-based AD framework that can perform close-loop autonomous driving in realistic simulators. To this end, (1) we bridge the gap between the language decisions and the vehicle control commands by standardizing the decision states according to the off-the-shelf motion planning module. (2) We employ a multi-modal LLM (MLLM) to model the behavior planning module of a module AD system, which uses driving rules, user commands, and inputs from various sensors (e.g., camera, lidar) as input and makes driving decisions and provide explanations; This model can plug-and-play in existing AD systems such as Apollo for close-loop driving. (3) We design an effective data engine to collect a dataset that includes decision state and corresponding explanation annotation for model training and evaluation. We conduct extensive experiments and show that our model achieves 76.1 driving score on the CARLA Town05 Long, and surpasses the Apollo baseline by 4.7 points under the same settings, demonstrating the effectiveness of our model. We hope this work can serve as a baseline for autonomous driving with LLMs.
@@ 1 Introduction
(a) Rule-Based Autonomous Driving System [3]
(b) End-to-End Autonomous Driving System [25, 27, 57]
@@
Figure 1: Comparison of different AD systems. (a) is the rule-based system constrained by the limitation of pre-defined and manual rules, (b) is a data-driven system restricted by the diversity of generated data, (c) we apply the capability of LLMs to the AD system and align the decision state with the planner to achieve closed-loop driving.
*** equal contribution, 🖂 corresponding author (daijifeng@tsinghua.
edu.cn)
@@
Autonomous driving (AD) has undergone significant advancements in recent years, evolving from traditional rule-based systems, which rely on a predefined set of rules informed by prior knowledge (see Figure 0(a)), to data-driven, end-to-end systems, as demonstrated in Figure 0(b). Despite their advancements, these systems have encountered limitations due to the constraints of expert knowledge or the diversity of training data. This has made it challenging for them to handle corner-case situations, even though human drivers might find them intuitive to deal with. In contrast to these traditional rule-based or data-driven AD planners, Large language models (LLMs) trained with web-scale text corpus, are equipped with extensive world knowledge, robust logical reasoning, and advanced cognitive capabilities. These features position them as potential planners in AD systems, providing a human-like approach to autonomous driving.
@@
Some recent studies [16, 72, 68, 39, 24, 13, 56] have been made to integrate LLMs into AD systems, focusing on generating language-based decisions in response to driving scenarios. However, these approaches have limitations when it comes to performing closed-loop driving in real-world environments or realistic simulators. This is because the outputs of LLMs are mainly linguistic and conceptual, which cannot be used for vehicle control. In traditional modular AD systems [3, 22, 21], the gap between high-level strategic goals and low-level operational actions is connected by a behavioral planning module, whose decision states can be easily transformed into vehicle control signals by follow-up motion planning and control. This motivates us to align the LLM with the decision state of the behavioral planning module, and further design an LLM-based close-loop AD system that can run on real-world environments or realistic simulators by using the aligned LLM for behavioral planning.
@@
Based on this point, we propose DriveMLM, the first LLM-based AD framework that can perform close-loop autonomous driving in realistic simulators. To achieve this, we have three key designs: (1) We investigate the decision states of the behavioral planning module of the well-developed Apollo system [3], and transform them into forms that can be easily processed by LLMs. (2) We develop a multi-modal LLM (MLLM) planner that can accept the current multi-modal inputs including multi-view images, LiDAR point clouds, traffic rules, system messages, and user instructions, and predict the decision state; (3) To obtain enough training data for behavioral planning state alignment, we manually collect 280 hours of driving data on CARLA, and convert them into decision state and corresponding explanation annotations by an efficient data engine. With these designs, we can obtain an MLLM planner that can make decisions based on the driving scenes and user requirements, and its decisions can be easily converted into vehicle control signals for closed-loop driving.
Our work has the following advantages: (1) Benefiting from the aligned decision states, our MLLM planner can be easily integrated with existing modular AD systems, such as Apollo, to achieve closed-loop driving without requiring any major changes or modifications. (2) By taking language instruction as input, our model can handle both user needs (e.g., overtaking a car) and high-level system messages (e.g., defining basic driving logic). This makes our model more flexible and adaptable to different driving situations and corner cases. (3) It can provide interpretability and explain different decisions. This enhances the transparency and trustworthiness of our model, as it can explain its actions and choices to the user.
In summary, the contribution of this work is three folds:
@@
(2) To implement this framework, we tailor a set of decision states with forms that can be easily processed by LLMs, design an MLLM planner for decision prediction, and develop a data engine that can effectively generate decision states and corresponding explanation annotation for model training and evaluation.
(3) To validate the effectiveness of our method, we not only evaluate our method on the closed-loop driving metrics including driving score (DS) and miles per intervention (MPI), but also use understanding metrics including accuracy, F1-measure for decision state, BLEU-4, CIDEr and METEOR for decision explanation to evaluate the driving understanding capability of our model. Notably, our method achieves 76.1 DS, 0.955 MPI results on CARLA Town05 Long, which is 4.7 points, 1.25 times better than Apollo. Moreover, we can change the decision of the MLLM planner by describing special requirements with language instructions such as yielding for ambulance or traffic rules, as shown in Figure 2.
@@ 2.1 Multi-Modal Large Language Models
The swift evolution of Large Language Models (LLMs) [53, 54, 7, 47, 46] has recently given rise to the emergence of multi-modal LLMs (MLLMs) [1, 26, 38, 37, 83, 17, 78, 12, 67, 51, 30, 2, 23, 33, 79, 71, 29, 80], which augment language models with the capacity to analyze and comprehend information from diverse modalities. Prominent instances of such advancements include GPT-4 [46], Flamingo [1], KOSMOS-1 [26], LLaVA series [38, 37], and MiniGPT-4 [83], as well as InstructBLIP [17]. These models have integrated visual instruction tuning methodologies to enhance the MLLMs’ ability to adhere to prescribed instructions. Furthermore, mPLUG-DocOwl [78] has broadened the document comprehension capabilities of MLLMs by incorporating digital document datasets. Concurrently, Shikra [12], VisionLLM [67], KOSMOS-2 [51], LISA [30], and Qwen-VL [2] have augmented MLLMs with visual grounding capabilities, empowering them to detect or segment objects in accordance with user prompts. The introduction of VideoChat [33] and VideoLLaMA [79] has ushered in the integration of video processing capabilities into LLMs. Additionally, NExT-GPT [71] has introduced a modality-switching instruction tuning technique for multi-modal prompt tuning, facilitating the handling of inputs and outputs in any combination of text, images, videos, and audio. ASM [29] and GPT4RoI [80] introduce region-level recognition and understanding capability into LLMs. These endeavors demonstrate the effectiveness and generalizability of LLMs, establishing a foundation for open-world tasks.
@@ 2.2 Intelligent Agents with Large Language Models
A burgeoning application of LLMs is their role in facilitating interaction and communication among intelligent agents (e.g., robots, virtual assistants, or game characters) and various entities, including humans, the environment, or even the intelligent agents themselves. Several API-based methods, including Visual ChatGPT [69], MM-REACT [77], HuggingGPT [59], InternGPT [40], ViperGPT [62], ControlLLM [41], and GPT4Tool [76] have attempted to integrate diverse modal APIs with LLMs to accomplish complex tasks in the open world, such as image editing, video processing, and audio synthesis. These methods allow language models to perform complex real-world tasks by following natural language instructions. In parallel, alternative research initiatives, such as Camel [31], AutoGPT [75], MetaGPT [24] and Smallville [50], investigate the utility of LLMs in the context of role-playing conversations or communication games. Additionally, within the domain of embodied AI, works such as PaLM-E [19], EmbodiedGPT [45], and the RT series [5, 6, 48] leverage LLMs to generate natural language actions, thereby controlling embodied agents proficient in executing navigation, manipulation, and interaction tasks within real or 3D environments. These works demonstrate the notable advancements achieved by LLMs in the realm of intelligent agent control.
@@ 2.3 Autonomous Driving Models
The development of autonomous driving (AD) models has accelerated rapidly in recent years, giving rise to many disruptive and groundbreaking technologies. Notably, the open-source frameworks, such as Apollo [3] and Autoware [22], have played pivotal roles by furnishing robust tools and resources, thereby facilitating the development of autonomous driving technology and contributing to its widespread adoption and progression. In terms of AD perception, BEV (Bird’s Eye View) [34, 73, 36, 61] and Occupancy Network [63, 60, 35] have become essential components of autonomous vehicles, helping them better understand the surrounding environment and make corresponding decisions. The decision-making process in conventional autonomous driving systems typically relies on finite state machines [14]. These systems often require the manual creation of numerous rules to determine the states and conditions for transitioning between them. However, considering the ever-changing nature of the world, this is usually laborious to design rules to cover all the scenarios for the real world. In recent years, end-to-end autonomous driving models have also made remarkable progress, such as UniAD [25], which adopts a novel end-to-end approach, directly integrating perception, prediction, and planning, avoiding information loss and efficiency issues in the traditional modular design method. Recently, open-sourced simulators [18, 66, 82] have been proposed to bridge the gap between model prediction and closed-loop control. Among them, CARLA [18], featuring comprehensive sensor simulations and realistic environments, is the most widely used benchmark for evaluating closed-loop performance by many state-of-the-art methods [27, 58, 57, 28, 15, 10, 11, 9].
Recent works [16, 72, 43, 68, 39, 13, 56] changes our perception by introducing LLM for driving planning, opening up a new direction for the autonomous driving field. As early explorations, some [68, 56] use ChatGPT and GPT-4 to predict driving decisions. Following works fine-tune LLM models to predict driving signal [13], trajectory [43] or designed decision space [39], conditioned only on language as input. DriveGPT4 [72] finetunes Multimodal LLM to predict control signal. However, DriveGPT4 is constrained by the input from a monocular camera, limiting its ability to construct comprehensive scene information. All LLM-based works above are not evaluated on realistic simulators in closed-loop driving, because either linguistic decisions of LLMs are hard to transform to actually reliable control signals, or the direct prediction of control signal by LLM remains a large gap to real-time closed-loop driving.
@@ 3.1 System Overview
Figure 3: DriveMLM framework consists of three parts: (1) behavioral planning states alignment, which aligns linguistic output to executable decisions for vehicle control. (2) MLLM planner, consisting of multi-model tokenizer and MLLM decoder. It transforms multi-modality sensor input $I,L$, system message $M$, and user instructions $U$ to drive explanation and aligned decisions. (3) Efficient data collection strategy, which generates rich driving explanations and aligned decisions at low cost. SPT stands for sparse pyramid transformer.
@@
The DriveMLM framework integrates the world knowledge and reasoning capabilities of large language models (LLMs) into an autonomous driving (AD) system, achieving closed-loop driving in realistic simulators. As illustrated in Figure 3, this framework has three key designs: (1) Behavioral Planning States Alignment. This part aligns LLM’s linguistic decision outputs with the behavioral planning module of a well-established modular AD system like Apollo [3]. In this way, the output of LLM can be easily transformed into vehicle control signals. (2) MLLM Planner. It is a combination of a multi-modal tokenizer and a multi-modal LLM (MLLM) decoder. The multi-modal tokenizer transforms diverse inputs like multi-view images, LiDAR, traffic rules, and user requirements into unified tokens, and the MLLM decoder makes decisions based on the unified tokens. (3) Efficient Data Collection Strategy. It introduces a tailored data collection method for LLM-based autonomous driving, ensuring a comprehensive dataset encompassing decision states, decision explanations, and user commands.
@@
During inference, the DriveMLM framework leverages multi-modal data to make driving decisions. These data include: multi-view images $I\!\in\!\mathbb{R}^{T\times N_{I}\times H\times W\times 3}$, where $T$ denotes the time length, $N_{I}$ indicates the number of views, and $H$ and $W$ denotes the height and width of images. The point clouds $L\!\in\!\mathbb{R}^{K\times 4}$ from LiDAR point clouds, with $K$ representing the number of points. System message $M\!\in\!\mathbb{R}^{N_{M}}$, with $N_{M}$ representing the number of system message tokens. The system message is the gathering of task definition, traffic rules, and decision state definition. User instructions $U\!\in\!\mathbb{R}^{N_{U}}$, where $N_{U}$ stands for the number of user instruction tokens. These inputs undergo tokenization through a multi-modal tokenizer, resulting in: $X_{I}\!\in\!\mathbb{R}^{N_{I}\times N_{Q}\times D}$, $X_{L}\!\in\!\mathbb{R}^{N_{Q}\times D}$, $X_{M}\!\in\!\mathbb{R}^{N_{M}\times D}$, $X_{U}\!\in\!\mathbb{R}^{N_{U}\times D}$, which represent the tokens embedding of multi-view images, LiDAR point clouds, traffic rules, and user instructions, respectively. Here, $N_{Q}$ denotes the output token number which is decided by the number of queries of QFormer [32], and each token embedding is with $D$ dimension. Next, these tokens are inputted into the MLLM decoder, which generates the decision state token $S$ along with a corresponding explanation $E$. Finally, the decision state $S$ is inputted into a motion planning and control module. This module computes the final trajectory for vehicle control.

### v3

# 8Beijing National Research Center for Information Science and Technology, Beijing, 100084, China Full list of author information is available at the end of the article Equal contributors

@@ Abstract
Large language models (LLMs) have opened up new possibilities for intelligent agents, endowing them with human-like thinking and cognitive abilities. In this work, we delve into the potential of large language models (LLMs) in autonomous driving (AD). We introduce DriveMLM, an LLM-based AD framework that can perform close-loop autonomous driving in realistic simulators. To this end, (1) we bridge the gap between the language decisions and the vehicle control commands by standardizing the decision states according to the off-the-shelf motion planning module. (2) We employ a multimodal LLM (MLLM) to model the behavior planning module of a module AD system, which uses driving rules, user commands, and inputs from various sensors (e.g., camera, lidar) as input and makes driving decisions and provide explanations; This model can plug-and-play in existing AD systems such as Autopilot and Apollo for close-loop driving. (3) We design an effective data engine to collect a dataset that includes decision state and corresponding explanation annotation for model training and evaluation. We conduct extensive experiments and show that replacing the decision-making modules of the Autopilot and Apollo with DriveMLM resulted in significant improvements of 3.2 and 4.7 points on the CARLA Town05 Long respectively, demonstrating the effectiveness of our model. We hope this work can serve as a baseline for autonomous driving with LLMs.
journal: Technical Report
@@ 1 Introduction
(a) Rule-Based Autonomous Driving System apollo_auto
(b) End-to-End Autonomous Driving System hu2023planning ; jia2023think ; shao2023safety
@@
Figure 1: Comparison of different AD systems. (a) is rule-based with manually defined rules, (b) is data-driven but lacks diversity in training data, and (c) integrates LLM capabilities with aligned decision states for closed-loop planning.
@@
Autonomous driving (AD) has undergone significant advancements in recent years, evolving from traditional rule-based systems, which rely on a predefined set of rules informed by prior knowledge (see Fig. 1(a)), to data-driven, end-to-end systems, as demonstrated in Fig. 1(b). Despite their advancements, existing end-to-end models are limited to learning from autonomous driving scenario data, lacking sufficient understanding of the real world, hindering the models’ ability to handle corner-case scenarios or those requiring complex reasoning. In contrast to traditional AD systems, Large language models (LLMs) trained with web-scale text corpus, are equipped with extensive world knowledge, robust logical reasoning, and advanced cognitive capabilities. These features position them as potential planners in AD systems, providing a human-like approach to autonomous driving.
@@
Some recent studies drivelm2023 ; xu2023drivegpt4 ; wen2023dilu ; liu2023mtd ; hong2023metagpt ; chen2023driving ; sha2023languagempc have been made to integrate LLMs into AD systems, focusing on generating language-based decisions in response to driving scenarios. However, these approaches have limitations when it comes to performing closed-loop driving in real-world environments or realistic simulators. This is because the outputs of LLMs are mainly linguistic and conceptual, which cannot be used for vehicle control. In traditional modular AD systems apollo_auto ; autoware ; fontana2021openpilot , the gap between high-level strategic goals and low-level operational actions is connected by a behavioral planning module, whose decision states can be easily transformed into vehicle control signals by follow-up motion planning and control. This motivates us to align the LLM with the decision state of the behavioral planning module, and further design an LLM-based close-loop AD system that can run on real-world environments or realistic simulators by using the aligned LLM for behavioral planning.
@@
Based on this point, we propose DriveMLM, an LLM-based AD framework that can perform close-loop autonomous driving in realistic simulators. To achieve this, we have three key designs: (1) We investigate the decision states of the behavioral planning module of the well-developed Apollo system apollo_auto , and transform them into forms that can be easily processed by LLMs. (2) We develop a multi-modal LLM (MLLM) planner that can accept the current multi-modal inputs including multi-view images, LiDAR point clouds, traffic rules, system messages, and user instructions, and predict the decision state; (3) To obtain enough training data for behavioral planning state alignment, we manually collect 280 hours of driving data on CARLA dosovitskiy2017carla , and convert them into decision state and corresponding explanation annotations by an efficient data engine. With these designs, we can obtain an MLLM planner that can make decisions based on the driving scenes and user requirements, and its decisions can be easily converted into vehicle control signals for closed-loop driving.
Our work offers several advantages, including: (1) Plug-and-Play. Our MLLM planner’s alignment with decision states allows seamless integration with existing modular autonomous driving (AD) systems like Apollo apollo_auto and AutoPilot tesla_autopilot . This integration enables closed-loop driving without significant alterations or adjustments to the existing systems. (2) Interactive. Leveraging the common knowledge and logical reasoning abilities of LLMs and taking language instructions as input, our model can comprehend high-level system messages, such as basic driving logic and complex user instructions. This makes our model more flexible and adaptable to diverse driving scenarios and corner cases.
In summary, our main contributions are three folds:
@@
(2) To implement this framework, we tailor a set of decision states with forms that can be easily processed by LLMs, design an MLLM planner for decision prediction. In addition, we also develop a data engine that can effectively generate decision states and the corresponding explanation annotation for model training and evaluation.
(3) To validate the effectiveness of our method, we not only evaluate our method on the closed-loop driving metrics including driving score (DS) and miles per intervention (MPI), but also use understanding metrics including accuracy, F1-measure for decision state, BLEU-4, CIDEr and METEOR for decision explanation to evaluate the driving understanding capability of our model. Notably, our method achieves 76.1 DS, 0.955 MPI results on CARLA Town05 Long, which is 4.7 points, 1.25 times better than Apollo apollo_auto . Moreover, our method demonstrates strong decision generalization capabilities when there are special requirements such as language instructions to hail an ambulance or obey traffic rules, as shown in Fig. 2.
@@ 2.1 Multi-modal Large Language Models
The swift evolution of Large Language Models (LLMs) radford2018improving ; radford2019language ; brown2020language ; ouyang2022training ; gpt4 has recently given rise to the emergence of multi-modal LLMs (MLLMs) alayrac2022flamingo ; huang2023language ; liu2023visual ; liu2023improved ; zhu2023minigpt ; instructblip ; ye2023mplug ; chen2023shikra ; wang2023visionllm ; peng2023kosmos ; lai2023lisa ; bai2023qwen ; gao2023llama ; li2023videochat ; zhang2023video ; wu2023next ; junqing2023never ; zhang2023gpt4roi , which augment language models with the capacity to analyze and comprehend information from diverse modalities. Prominent instances of such advancements include GPT-4 gpt4 , Flamingo alayrac2022flamingo , KOSMOS-1 huang2023language , LLaVA series liu2023visual ; liu2023improved , and MiniGPT-4 zhu2023minigpt , as well as InstructBLIP instructblip . These models have integrated visual instruction tuning methodologies to enhance the MLLMs’ ability to adhere to prescribed instructions.
Furthermore, mPLUG-DocOwl ye2023mplug has broadened the document comprehension capabilities of MLLMs by incorporating digital document datasets. Concurrently, Shikra chen2023shikra , VisionLLM wang2023visionllm , KOSMOS-2 peng2023kosmos , LISA lai2023lisa , and Qwen-VL bai2023qwen have augmented MLLMs with visual grounding capabilities, empowering them to detect or segment objects in accordance with user prompts. LLaVA-NeXt liu2023visual and InternVL-2 chen2024internvl ; chen2024far leavage dynamic image splitting to handle high-resolution images. Eagle-2 li2024eagle2fasterinferencelanguage empolies different vision encoders including both SigLIP zhai2023sigmoid and ConvNeXt liu2022convneXt that enhances the visual features. The introduction of VideoChat li2023videochat and VideoLLaMA zhang2023video has ushered in the integration of video processing capabilities into LLMs. Additionally, NExT-GPT wu2023next has introduced a modality-switching instruction tuning technique for multi-modal prompt tuning, facilitating the handling of inputs and outputs in any combination of text, images, videos, and audio. ASM junqing2023never and GPT4RoI zhang2023gpt4roi introduce region-level recognition and understanding capability into LLMs. These endeavors demonstrate the effectiveness and generalizability of LLMs, establishing a foundation for open-world tasks.
@@ 2.2 Intelligent Agents with Large Language Models
A burgeoning application of LLMs is their role in facilitating interaction and communication among intelligent agents (e.g., robots, virtual assistants, or game characters) and various entities, including humans, the environment, or even the intelligent agents themselves. Several API-based methods, including Visual ChatGPT wu2023visual , MM-REACT yang2023mm , HuggingGPT shen2023hugginggpt , InternGPT liu2023internchat , ViperGPT suris2023vipergpt , ControlLLM liu2023controlllm , and GPT4Tool yang2023gpt4tools have attempted to integrate diverse modal APIs with LLMs to accomplish complex tasks in the open world, such as image editing, video processing, and audio synthesis. These methods allow language models to perform complex real-world tasks by following natural language instructions. In parallel, alternative research initiatives, such as Camel li2023camel , AutoGPT yang2023auto , MetaGPT hong2023metagpt and Smallville park2023generative , investigate the utility of LLMs in the context of role-playing conversations or communication games. Additionally, within the domain of embodied AI, works such as PaLM-E driess2023palm , EmbodiedGPT mu2023embodiedgpt , and the RT series brohan2022rt ; brohan2023rt ; padalkar2023open leverage LLMs to generate natural language actions, thereby controlling embodied agents proficient in executing navigation, manipulation, and interaction tasks within real or 3D environments. These works demonstrate the notable advancements achieved by LLMs in the realm of intelligent agent control.
@@ 2.3 Autonomous Driving Models
The development of autonomous driving (AD) models has accelerated rapidly in recent years, giving rise to many disruptive and groundbreaking technologies. Notably, the open-source frameworks, such as Apollo apollo_auto and Autoware autoware , have played pivotal roles by furnishing robust tools and resources, thereby facilitating the development of autonomous driving technology and contributing to its widespread adoption and progression. In terms of AD perception, BEV (Bird’s Eye View) li2022bevformer ; yang2023bevformer ; liang2022bevfusion ; singh2023surround and Occupancy Network tong2023scene ; shi2023grid ; li2023fb have become essential components of autonomous vehicles, helping them better understand the surrounding environment and make corresponding decisions.
The decision-making process in conventional autonomous driving systems typically relies on finite state machines chen2019autonomous . These systems often require the manual creation of numerous rules to determine the states and conditions for transitioning between them. However, considering the ever-changing nature of the world, this is usually laborious to design rules to cover all the scenarios for the real world. In recent years, end-to-end autonomous driving models have also made remarkable progress, such as UniAD hu2023planning , which adopts a novel end-to-end approach, directly integrating perception, prediction, and planning, avoiding information loss and efficiency issues in the traditional modular design method.
Recently, open-sourced simulators dosovitskiy2017carla ; vinitsky2022nocturne ; zhou2020smarts have been proposed to bridge the gap between model prediction and closed-loop control. Among them, CARLA dosovitskiy2017carla , featuring comprehensive sensor simulations and realistic environments, is the most widely used benchmark for evaluating closed-loop performance by many state-of-the-art methods jia2023think ; shao2023reasonnet ; shao2023safety ; jiang2023vad ; chitta2022transfuser ; chen2020learning ; chen2021learning ; chen2022learning .
Recent works drivelm2023 ; xu2023drivegpt4 ; mao2023gpt ; wen2023dilu ; liu2023mtd ; chen2023driving ; sha2023languagempc changes our perception by introducing LLM for driving planning, opening up a new direction for the autonomous driving field.As early explorations, some wen2023dilu ; sha2023languagempc use ChatGPT and GPT-4 to predict driving decisions. Following works fine-tune LLM models to predict driving signal chen2023driving , trajectory mao2023gpt or designed decision space liu2023mtd , conditioned only on language as input. DriveGPT4 xu2023drivegpt4 finetunes Multimodal LLM to predict control signal. However, DriveGPT4 is constrained by the input from a monocular camera, limiting its ability to construct comprehensive scene information.
All LLM-based works above are not evaluated on realistic simulators in closed-loop driving, because either linguistic decisions of LLMs are hard to transform to actually reliable control signals, or the direct prediction of control signal by LLM remains a large gap to real-time closed-loop driving.
@@ 3.1 System Overview
Figure 3: DriveMLM framework consists of three parts: (1) behavioral planning states alignment, which aligns linguistic output to executable decisions for vehicle control. (2) MLLM planner, consisting of multi-modal tokenizer and MLLM decoder. It transforms multi-modality sensor input $I,L$, system message $M$, and user instructions $U$ to drive explanation and aligned decisions. (3) Efficient data collection strategy, which generates rich driving explanations and aligned decisions at low cost. SST stands for single-stride sparse transformer.
@@
The DriveMLM framework integrates the world knowledge and reasoning capabilities of large language models (LLMs) into an autonomous driving (AD) system, achieving closed-loop driving in realistic simulators. As illustrated in Fig. 3, this framework has three key designs: (1) Behavioral Planning States Alignment. This part aligns LLM’s linguistic decision outputs with the behavioral planning module of a well-established modular AD system like autopilot tesla_autopilot . In this way, the output of LLM can be easily transformed into vehicle control signals. (2) MLLM Planner. It is a combination of a multi-modal tokenizer and a multi-modal LLM (MLLM) decoder. The multi-modal tokenizer transforms diverse inputs like multi-view images, LiDAR, traffic rules, and user requirements into unified tokens, and the MLLM decoder makes decisions based on the unified tokens. (3) Efficient Data Collection Strategy. It introduces a tailored data collection method for LLM-based autonomous driving, ensuring a comprehensive dataset encompassing decision states, decision explanations, and user commands.
@@
During inference, the DriveMLM framework leverages multi-modal data to make driving decisions. These data include: multi-view images $I\!\in\!\mathbb{R}^{T\times N_{I}\times H\times W\times 3}$, where $T$ denotes the time length, $N_{I}$ indicates the number of views, and $H$ and $W$ denotes the height and width of images. The point clouds $L\!\in\!\mathbb{R}^{K\times 4}$ from LiDAR point clouds, with $K$ representing the number of points. System message is denoted by $M\!\in\!\mathbb{R}^{N_{M}}$, and $N_{M}$ represents the number of system message tokens. The system message is the gathering of task definition, traffic rules, and decision state definition. User instructions $U\!\in\!\mathbb{R}^{N_{U}}$, where $N_{U}$ stands for the number of user instruction tokens. These inputs undergo tokenization through a multi-modal tokenizer, resulting in: $X_{I}\!\in\!\mathbb{R}^{N_{I}\times N_{Q}\times D}$, $X_{L}\!\in\!\mathbb{R}^{1\times D}$, $X_{M}\!\in\!\mathbb{R}^{N_{M}\times D}$, $X_{U}\!\in\!\mathbb{R}^{N_{U}\times D}$, which represent the tokens embedding of multi-view images, LiDAR point clouds, traffic rules, and user instructions, respectively. Here, $N_{Q}$ denotes the number of output image token which is decided by the number of features in the last Transformer layer of the pre-trained CLIP visual encoder, and each token embedding is with $D$ dimension. Next, these tokens are inputted into the MLLM decoder, which generates the decision state token $S$ along with a corresponding explanation $E$. Finally, the decision state $S$ is inputted into a motion planning and control module. This module computes the final trajectory for vehicle control.